# Boosting Graph Pooling with Persistent Homology

### v1

@@ E.3 Visualization of coarsened graphs without preserving topology
We present some coarsened graphs that do not preserve topology (DiffPool-TIP-NL) in Fig. 9. These graphs contribute equally to the objective in the graph classification task, but their topologies are different. A similar observation was made by Mesquita et al. (2020), who found that randomly generated graphs show equivalent performance. In DiffPool-TIP-NL, other topology-related modules in TIP are preserved, allowing some topological information to be injected into the three results shown in Fig. 9. Guided by the $\mathcal{L}_{topo}$, DiffPool-TIP tends to select the results that are most similar to the original graph among all the options. Experimental results in Fig. 5 demonstrate that this type of topology is superior and leads to better performance on downstream tasks.
@@
E.4 Ablation Study
@@
DMoNPool-TIP 79.68 $\pm$ 1.38 79.73 $\pm$ 3.66 45.84 $\pm$ 5.32 74.25 $\pm$ 3.06

### v2

@@ E.3 Visualization of coarsened graphs without preserving topology
We present some coarsened graphs that do not preserve topology (DiffPool-TIP-NL) in Fig. 9. These graphs contribute equally to the objective in the graph classification task, but their topologies are different. A similar observation was made by [29], who found that randomly generated graphs show equivalent performance. In DiffPool-TIP-NL, other topology-related modules in TIP are preserved, allowing some topological information to be injected into the three results shown in Fig. 9. Guided by the $\mathcal{L}_{topo}$, DiffPool-TIP tends to select the results that are most similar to the original graph among all the options. Experimental results in Fig. 5 demonstrate that this type of topology is superior and leads to better performance on downstream tasks.
@@
E.4 Topology relevant experiments
To further demonstrate that our proposed method can effectively capture the topological features in graphs, we design an experiment where the topological structure of the graph is highly relevant. We generate a synthetic dataset named Cycles, comprising two balanced 2-class sets of 1000 graphs each. This dataset consists of either a single large cycle (class 0) or two connected large cycles (class 1), resembling digital numbers “0” and “8”, respectively. The distinguishing factor between the classes lies in the presence of cycles, highlighting the significance of the graph’s topological structure in classification. The node numbers range from 10 to 20, with 3-dimensional random node features generated. For model configuration, we uniformly use 1-GCN plus 1-pooling layer. The evaluation criteria remain consistent with those outlined in our paper. The experimental results in Table 6 demonstrate the effectiveness of TIP in leveraging topological features to significantly outperform the comparable pooling methods.
Table 6: Classification results on Cycles dataset
Test Accuracy
DiffPool 54.3 $\pm$ 1.1
DiffPool-TIP 65.1 $\pm$ 2.7
MinCutPool 54.2 $\pm$ 2.6
MinCutPool-TIP 65.0 $\pm$ 2.9
DMoNPool 55.0 $\pm$ 2.7
DMoNPool-TIP 68.7 $\pm$ 2.7
E.5 Ablation study
@@
Further, we provide an ablation study of our topological loss term by replacing it with the Wasserstein distance. While the Wasserstein distance is a powerful metric for comparing persistence diagrams, its computation can be computationally intensive, particularly when dealing with high-dimensional vectorized representations. Therefore, it significantly increases our training time in practice. We denote the variant of using Wasserstein distance as “TIP-W”. Here we present the ablation study results on two datasets. We can observe that TIP-W has competitive performance compared with the full version TIP (with our proposed loss term), and outperforms the variant TIP-NL (no loss term). Initially, we design our $\mathcal{L}_{topo}$ to avoid the high computational complexity of Wasserstein distance, but we are suprised to find that TIP also marginally outperforms TIP-W in numerous instances, potentially attributed to the efficacy of feature transformation and high-order statistical features. These elements serve as a feature augmentation mechanism to enhance the persistence diagrams.
@@
DMoNPool-TIP-W 79.48 $\pm$ 1.50 79.70 $\pm$ 2.95 45.45 $\pm$ 1.34 74.35 $\pm$ 2.91
DMoNPool-TIP 79.68 $\pm$ 1.38 79.73 $\pm$ 3.66 45.84 $\pm$ 5.32 74.00 $\pm$ 2.93
E.6 Evaluation of expressive power
The growing interest in the expressive capability of graph pooling has been prominent in recent studies [3]. A graph pooling model based on GNNs is deemed more effective as it can differentiate a larger set of non-isomorphic graphs by producing unique representations for each. Graph pooling integrated with appropriately designed message-passing layers proves to be as competent as the WL test in distinguishing graphs. Understanding the expressive capacity of graph pooling aids in selecting between existing pooling operators or crafting novel ones. Furthermore, to empirically assess the expressive capacity of our proposed approach, TIP, we conduct experiments on the EXPWL1 dataset following the experimental setup detailed in [3]. Each graph pair $\left(\mathcal{G}_{i},\mathcal{H}_{i}\right)$ in EXPWL1 consists of two non-isomorphic graphs distinguishable by a WL test, which encode formulas with opposite SAT outcomes. Therefore, any GNN that has an expressive power equal to the WL test can distinguish them and achieve approximately $100\%$ classification accuracy on the dataset. The classification outcomes on the EXPWL1 dataset are shown in Table 8, which reveal the notable improvement in the expressive capacity of graph pooling achieved through our proposed method in empirical evaluations.
Table 8: Classification results on EXPWL1 dataset.
Pooling Test Accuracy
DiffPool 97.0 $\pm$ 2.4
DiffPool-TIP 99.3 $\pm$ 0.5
MinCutPool 98.8 $\pm$ 0.4
MinCutPool-TIP 99.9 $\pm$ 0.1
DMoNPool 99.0 $\pm$ 0.7
DMoNPool-TIP 99.7 $\pm$ 0.1
Appendix F Limitations
The potential limitation of our study is the heavy reliance of the proposed method on circular structures within graphs, potentially hindering its efficacy on tree-like graphs.